# ReGNet: Reciprocal Space-Aware Long-Range Modeling for Crystalline Property Prediction

### v2

@@ Abstract
Predicting properties of crystals from their structures is a fundamental yet challenging task in materials science. Unlike molecules, crystal structures exhibit infinite periodic arrangements of atoms, requiring methods capable of capturing both local and global information effectively. However, most current works fall short of capturing long-range interactions within periodic structures. To address this limitation, we leverage reciprocal space to efficiently encode long-range interactions with learnable filters within Fourier transforms. We introduce Reciprocal Geometry Network (ReGNet), a novel architecture that integrates geometric GNNs and reciprocal blocks to model short-range and long-range interactions, respectively. Experimental results on JARVIS, Materials Project, and MatBench demonstrate that ReGNet achieves state-of-the-art predictive accuracy across a range of crystal property prediction tasks. Additionally, we explore a model extension that employs the mixture-of-experts for multi-property prediction with promising results and high computational efficiency. These findings highlight the potential of our model as a scalable and accurate solution for crystal property prediction. The code will be released upon paper acceptance.
@@ 1 Introduction
AI-driven methods for accelerated material property prediction (Choudhary et al., 2022) have achieved notable success in materials sciences. As the demand for novel materials grows, traditional methods like wet lab experiments are costly and time-intensive. Computational methods such as density functional theory (DFT) offer advances but remain limited by high computational costs, hindering scalable exploration of the vast materials space (Jones, 2015). To overcome these challenges, recent studies have adopted machine learning (ML) models in material property prediction and accelerating discovery (Chen et al., 2019; Yan et al., 2022; Zeni et al., 2023; Lin et al., 2023; Yan et al., 2024).
Crystals are solid materials extensively utilized across various scientific and industrial domains, including semiconductors, photonics, catalysis, and pharmaceuticals (Ashcroft & Mermin, 2001). Unlike molecules, crystal structures possess a unique and essential structural feature called periodicity, in which atoms are arranged in a highly ordered, infinitely repeating pattern that extends in three-dimensional (3D) space, as illustrated in Figure 1. This periodic arrangement, referred to as the crystal lattice, fundamentally shapes the material’s distinct physical and chemical properties. Recently, graph neural network (GNN)-based methods have been developed to predict these properties by constructing graphs connecting atoms within a small, pre-defined distance threshold (Xie & Grossman, 2018; Schütt et al., 2018; Chen et al., 2019; Louis et al., 2020; Choudhary & DeCost, 2021). However, these methods fail to explicitly capture long-range interactions between distant atoms.
A key challenge in predicting crystal properties is effectively encoding long-range atomic interactions within periodic structures. The previous GNN-based method, PotNet (Lin et al., 2023), uses fixed-form physical equations to compute infinite potential summations as edge features. However, this method incurs an additional $\mathcal{O}(n^{2})$ computational cost, making it inefficient for complex material systems with large $n$, where $n$ is the number of atoms in the unit cell. Furthermore, material interactions are inherently complicated and cannot be adequately captured by a direct summation of a few predefined potentials.
Figure 1: Illustration of the relationship between real space (left) and reciprocal space (right) in crystals. Long-range interactions can be computed through the Fourier series expansion. See Section 2.2 for details.
In this work, we utilize reciprocal space to efficiently encode long-range interactions using learnable continuous filters with Fourier transforms in crystalline materials. Reciprocal space provides a mathematical framework to describe the periodic structure of crystals, enabling the analysis and prediction of material properties influenced by periodicity Economou (2010). While long-range interactions decay slowly with distance in real space, their Fourier transform decays rapidly, enabling efficient computations. Despite its fundamental importance, reciprocal space remains largely overlooked in crystal representation learning. Thus, integrating reciprocal space information into such models presents a novel opportunity to capture global periodicity and interactions, with the potential to substantially improve the accuracy of predictions.
Previous work, EwaldMP (Kosmala et al., 2023), encodes molecular long-range interactions using Fourier transforms, but only for molecular systems and relies on a predefined supercell grid size as hyperparameters, which can disrupt unit cell symmetry. In contrast, our method leverages reciprocal lattice vectors and atomic fractional coordinates, thus eliminating dependence on predefined supercell grids and enhancing symmetry preservation for more accurate modeling. Crystalformer (Taniai et al., 2024) introduces reciprocal space via fixed Fourier distance features to enable infinite range attention, where it is non-trainable and appears only as an optional variant in the appendix. In contrast, we utilize message passing with learnable continuous filters, parameterized by neural networks, to enable more precise representations of long-range interactions in complex crystals. In brief, this paper makes the following contributions.
  • $\bullet$
    Reciprocal space. We introduce the first end-to-end framework that incorporates reciprocal space with trainable Fourier-space filters into message passing for crystalline materials, enabling reciprocal space informed modeling of long-range interactions.
  • $\bullet$
    Architecture design. We introduce a new model, Reciprocal Geometry Network (ReGNet). It consists of multiple message-passing blocks, where each block comprises a geometric GNN for short-range interactions and a novel reciprocal block for long-range interactions.
  • $\bullet$
    Superior empirical performance. Extensive experiments show that ReGNet achieves state-of-the-art performance compared to various methods across common material property prediction datasets, JARVIS, the Material Project, and MatBench datasets.
@@ 2.1 Periodic Crystal Structure Representation
A crystalline material can be represented as the infinite periodic arrangement of atoms in 3D space, with the smallest repeating structural unit referred to as the unit cell. The unit cell defines the lattice symmetry and dictates the overall periodicity and physical properties of the material. The structure of crystals is commonly represented using two coordinate systems: the Cartesian coordinate system and the fractional coordinate system.
@@
Mathematically, a material can be represented as $\mathbf{M}=(\mathbf{A},\mathbf{P},\mathbf{L})$. $\mathbf{A}=[\bm{a}_{1},\bm{a}_{2},\cdots,\bm{a}_{n}]^{\top}\in\mathbb{R}^{n\times h}$ denotes the atom feature matrix for $n$ atoms within a unit cell, where $\bm{a}_{i}\in\mathbb{R}^{h}$ represents the $h$-dimensional features of atom $i$ in the unit cell. $\textbf{P}=[\bm{p}_{1},\bm{p}_{2},\cdots,\bm{p}_{n}]^{\top}\in\mathbb{R}^{n\times 3}$ is the position matrix, where $\bm{p}_{i}\in\mathbb{R}^{3}$ is the Cartesian coordinate of atom $i$ inside the unit cell. In crystallography, a lattice is a periodic arrangement of points that defines a crystal’s translational symmetry. As shown in Figure 1, the lattice matrix $\textbf{L}=[\bm{\ell}_{1},\bm{\ell}_{2},\bm{\ell}_{3}]^{\top}\in\mathbb{R}^{3\times 3}$ consists of three translation lattice vectors that determine the shape and periodicity of the unit cell. The periodic repetition of the unit cell is described by integer multiples of the lattice vectors, ensuring translational invariance. Specifically, the infinite crystal structure can be expressed as:
[…]
where the integers $n_{1},n_{2},n_{3}$ define a 3D translation with $\bm{\ell}_{1},\bm{\ell}_{2},\bm{\ell}_{3}$. $\hat{\mathbf{A}}=\{\hat{\bm{a}_{i}}|\hat{\bm{a}_{i}}=\bm{a}_{i},i\in\mathbb{Z},1\leq i\leq n\}$ is the atom feature in repeated unit cells, which remains unchanged under periodic translation.
Figure 2: Overall framework architecture, consisting of three components: (a) Model architecture, (b) ReGNet block, and (c) Decoder. The model uses stacked ReGNet blocks, which integrate local features via message passing and global features via reciprocal space convolution. The decoder outputs property predictions for single-task and multi-task settings.
@@
Fractional coordinates capture the relative positions of atoms within the unit cell and offer significant advantages for periodic materials compared to Cartesian coordinates. Specifically, the fractional coordinate uses the lattice matrix $\textbf{L}\in\mathbb{R}^{3\times 3}$ as the basis, where atomic positions are described by fractional coordinate vectors $\bm{f_{i}}=[f_{1},f_{2},f_{3}]^{\top}\in[0,1)^{3}$. The corresponding Cartesian coordinates are obtained as $\bm{p}_{i}=\bm{L}\bm{f_{i}}$, where $\bm{p}_{i}=f_{1}\bm{\ell}_{1}+f_{2}\bm{\ell}_{2}+f_{3}\bm{\ell}_{3}$. The representation of crystal $\mathbf{M}$ generalizes to $\mathbf{M}=(\mathbf{A},\mathbf{F},\mathbf{L})$, where $\mathbf{F}=[\bm{f}_{1},\cdots,\bm{f}_{n}]^{\top}\in[0,1)^{n\times 3}$ denotes the fractional coordinates of all atoms in the unit cell. Fractional coordinates inherently align with lattice periodicity, ensuring that calculations respect periodic boundary conditions. Moreover, they remain invariant under space group operations, such as translations and rotations, thereby preserving structural relationships. These properties make fractional coordinates a more efficient and symmetry-aware representation for periodic materials compared to Cartesian coordinates.
@@ 2.2 Reciprocal Space in Crystalline Materials
The periodic arrangement of atoms in crystalline materials fundamentally influences their physical properties by establishing a repeating structure that extends over long distances. Reciprocal space provides a mathematical framework to describe the periodic structure of crystals, enabling the analysis and prediction of material properties influenced by periodicityAshcroft & Mermin (2001); Gross & Marx (2014); Economou (2010). Reciprocal space consists of wave vectors $\bm{k}$ spanned by reciprocal lattice vectors $\bm{b_{1}},\bm{b_{2}},\bm{b_{3}}$, which are derived from lattice vectors in real space $\bm{\ell}_{1},\bm{\ell}_{2},\bm{\ell}_{3}$, with relationship satisfying $\bm{b_{i}}\cdot\bm{\ell_{j}}=2\pi\delta_{ij}$, where $\delta_{ij}$ is the Kronecker delta. Each vector in reciprocal space represents a spatial frequency that encodes crystal periodicity along specific directions. More details can be found in Appendix A.
To establish the reciprocal space representation, we employ the Fourier transform, which expresses the periodic function $f$ via Fourier series expansion:
@@
where $\hat{f}(\bm{k})$ represents the Fourier coefficients, encoding the contributions of wave vectors $\bm{k}$ at atom $j$. The corresponding inverse transform, which computes $\hat{f}(\bm{k})$ from the real-space representation $f(\bm{p}_{j})$, is given by:
@@ 3 Method
This work focuses on predicting the thermodynamic, electronic, and mechanical properties of crystalline materials through a novel approach that explicitly incorporates long-range interactions from reciprocal space. The architectures of ReGNet and its multi-task variant, ReGNet-MT, are illustrated in Figure 2. Unlike Lin et al. (2023), which employs fixed physical potentials as physics-informed edge features, our approach employs continuous filters with Fourier transforms, utilizing learnable parameters to encode reciprocal space representations while preserving the periodic symmetry of crystalline systems. Inspired by the molecular graph framework in Kosmala et al. (2023), we propose the ReciprocalBlock, a specialized module designed to capture long-range interactions in crystalline materials. This module leverages reciprocal lattice vectors and atomic fractional coordinates, eliminating reliance on predefined supercell grid sizes. Therefore, our approach enhances symmetry and periodicity preservation in long-range representations. Besides, we incorporate GNN with our ReciprocalBlock to model short- and long-range information in materials. Lastly, we enable a variant multi-property prediction by leveraging mixture-of-experts (MoE) in the decoder.

### v3

# ReciNet: Reciprocal Space-Aware Long-Range Modeling for Crystalline Property Prediction

@@ Abstract
Predicting properties of crystals from their structures is a fundamental yet challenging task in materials science. Unlike molecules, crystal structures exhibit infinite periodic arrangements of atoms, requiring methods capable of capturing both local and global information effectively. However, current works fall short of capturing long-range interactions within periodic structures. To address this limitation, we leverage reciprocal space, the natural domain for periodic crystals, and construct a Fourier series representation from fractional coordinates and reciprocal lattice vectors with learnable filters. Building on this principle, we introduce the reciprocal space-based geometry network (ReciNet), a novel architecture that integrates geometric GNNs and reciprocal blocks to model short-range and long-range interactions, respectively. Experimental results on standard benchmarks JARVIS, Materials Project, and MatBench demonstrate that ReciNet achieves state-of-the-art predictive accuracy across a range of crystal property prediction tasks. Additionally, we explore a model extension to multi-property prediction with the mixture-of-experts, which demonstrates high computational efficiency and reveals positive transfer between correlated properties. These findings highlight the potential of our model as a scalable and accurate solution for crystal property prediction.
@@ 1 Introduction
AI-driven methods are revolutionizing materials science (Choudhary et al., 2022) by accelerating the discovery of novel materials that are critical for technological applications, including semiconductors, catalysts, and pharmaceuticals (Ashcroft & Mermin, 2001). In parallel, although traditional computational approaches, such as density functional theory (DFT), have provided significant advances, their high computational cost limits their scalability for exploring the vast materials space (Jones, 2015). To overcome these challenges, machine learning (ML) models, such as Graph Neural Networks (GNNs), have become prevailing for predicting material properties directly from their atomic structures (Chen et al., 2019; Louis et al., 2020; Zeni et al., 2023; Lin et al., 2023; Yan et al., 2024). However, these models suffer from a fundamental limitation that most approaches only capture local information, which introduces a locality bias. Message passing is typically restricted to atoms within a fixed cutoff radius, without capturing crystalline infinite, periodic atomic arrangement, and the associated long-range information, as illustrated in Figure˜1 (Xie & Grossman, 2018; Schütt et al., 2018; Yan et al., 2022; Choudhary & DeCost, 2021). This limitation is particularly problematic because many key material properties, such as electronic band structures and mechanical moduli, are strongly dependent on this long-range information, creating a critical mismatch between conventional GNN paradigms and the underlying physics of crystalline solids.
The natural and principled framework for describing periodicity and long-range phenomena in solid-state physics is reciprocal space (Economou, 2010). By applying a Fourier transform, interactions that decay slowly in real space are represented compactly in reciprocal space, where only a few strong, low-frequency components dominate. This transformation makes modeling long-range interactions computationally tractable and provides a physically grounded strategy for overcoming the locality bias. However, designing an effective, end-to-end machine learning architecture that leverages reciprocal space for crystalline systems has remained a significant and largely unmet challenge.
A few early attempts have been made to integrate reciprocal space for crystalline property prediction, but these preliminary efforts fall short due to critical design limitations. One line of work models long-range interactions through fixed, non-trainable formulations. PotNet (Lin et al., 2023) encodes precomputed physical potentials derived from Ewald summation, while Crystalformer (Taniai et al., 2024) incorporates a dual-space attention mechanism based on analytical Fourier-Gaussian kernels. However, these methods rely on rigid, handcrafted functions that prevent the model from learning material-specific reciprocal space interactions. Consequently, they provide only a shallow integration of reciprocal space and fail to fully leverage the symmetry-aware structure intrinsic to periodic crystals. A separate line of work, while offering a learnable design, suffers from other fundamental shortcomings. EwaldMP (Kosmala et al., 2023) introduces an end-to-end approach with structure factor embeddings. However, it was developed primarily for molecular systems and relies on an artificial supercell grid. This formulation is fundamentally misaligned with the physics of crystalline solids, as it breaks the crystal’s inherent space group symmetries and introduces computationally costly grid-size hyperparameters. Together, these issues highlight a clear and critical gap: the lack of a crystal-native framework that simultaneously preserves intrinsic periodic symmetries and adaptively learns complex interactions in reciprocal space.
Figure 1: Schematic illustrations of capturing interactions in real space and reciprocal space. Note that we model 3D crystals, while we have a 2D illustration for simplicity. (a) An example crystal where each unit cell contains two atoms $i$ and $j$. Methods that rely solely on local information, such as CGCNN(Xie & Grossman, 2018), apply a fixed cutoff radius, thereby neglecting long-range interatomic interactions that are essential for accurately predicting crystal properties. (b) Relationship between real space (left) and reciprocal space (right). Reciprocal space enables efficient computation of long-range interactions via Fourier series expansion. See Section˜3.2 for additional background.
In this work, we bridge this gap by introducing the reciprocal space-based geometry network (ReciNet), a neural network architecture designed to capture long-range interactions in a crystal-native and system-adaptive manner. Our key contribution lies in the reciprocal block, a novel module that models long-range interactions by constructing a learnable Fourier series representation from fractional coordinates and reciprocal lattice vectors. This design inherently preserves crystal symmetries and adapts to diverse material systems without relying on fixed grids or analytical functions. Then, ReciNet integrates this reciprocal block with a local geometric GNN, forming a hybrid architecture that jointly models short-range chemical bonding and long-range lattice-scale effects. We further extend the framework to a multi-task variant, enabling more efficient and scalable multi-property prediction. Finally, extensive experiments show that ReciNet achieves state-of-the-art performance compared to various methods across common material property prediction datasets, JARVIS (Choudhary et al., 2020), the Materials Project (Chen et al., 2019), and MatBench datasets (Dunn et al., 2020). Especially, ReciNet achieves 13.7% lower MAE in bulk modulus prediction and 7.7% in band gap (MBJ) prediction compared with iComFormer. Lastly, we show the superiority of our method compared to prior methods that incorporate reciprocal space information.
2 Related Work
Crystal property prediction.
Crystal property prediction has been widely studied using both physics-based and deep-learning approaches. Traditional physics-based methods, such as Coulomb matrices (Rupp et al., 2012; Elton et al., 2018), effectively model ionic and metallic materials but lack generalization due to constraints like permutation invariance. Deep learning has introduced more flexible predictions by representing crystals as chemical formulas and using sequence models (Jha et al., 2018; Goodall & Lee, 2020; Wang et al., 2021). More recent methods leverage 3D geometric structures, modeling as 3D graphs in SchNet (Schütt et al., 2017) and DimeNet (Gasteiger et al., 2020). Key advances include CGCNN (Xie & Grossman, 2018), which utilizes multi-edge graphs to model periodic invariance. ALIGNN (Choudhary & DeCost, 2021) incorporates angle-based features, and Matformer (Yan et al., 2022) captures periodic patterns using self-connecting edges. Yan et al. (2024) introduces iComFormer with invariant descriptors and eComFormer with equivariant vectors. CrystalFramer (Ito et al., 2025) models SE(3)-invariant features using dynamic, atom-wise coordinate frames. Despite advancements, these approaches face significant limitations in accurately capturing periodic patterns and long-range interactions.
Long-range interaction modeling.
Modeling long-range interactions in periodic crystals remains a significant challenge. Early data-driven methods augmented local message-passing networks with simplified global information, such as predicting atomic charges (Unke et al., 2021) or integrating computed scalar-valued structure factors only at the final layer (Yu et al., 2022), which limits the rich interplay between local and global features. More physically-grounded approaches operate in reciprocal space, the natural domain for periodic systems (Ashcroft & Mermin, 2001). However, these models often suffer from the following critical limitations. PotNet (Lin et al., 2023) uses pre-computed potential fields derived from Ewald summation as static node features, while Crystalformer (Taniai et al., 2024) employs an attention mechanism based on fixed, analytical Fourier-Gaussian kernels. While physically motivated, these hand-crafted functions are universal across all materials. Moreover, the pre-computed potential in PotNet requires high $\mathcal{O}(n^{2})$ computational cost , where $n$ is the number of atoms in the unit cell. Another approach adopts a non-crystal-native formulation. EwaldMP (Kosmala et al., 2023), for example, computes learnable structure factor embeddings but then projects them onto an artificial supercell grid to perform reciprocal-space message passing. This grid-based approach is fundamentally misaligned with true crystal periodicity, as the arbitrary choice of grid size can break the crystal’s intrinsic space group symmetries and introduce computationally expensive hyperparameters. We include a direct comparison with reciprocal space methods in LABEL:{exp:reciprocal}.
@@ 3.1 Periodic Crystal Structure Representation
A crystalline material can be represented as the infinite periodic arrangement of atoms in 3D space, with the smallest repeating structural unit referred to as the unit cell (Ashcroft & Mermin, 2001). As the fundamental repeating unit, the unit cell contains the atomic arrangement and lattice vectors that establish a crystal’s periodicity and space group symmetry, which together dictate all of its material properties. The structure of crystals is commonly represented using two coordinate systems: the Cartesian coordinate system and the fractional coordinate system.
@@
Mathematically, a material can be represented as $\mathbf{M}=(\mathbf{A},\mathbf{P},\mathbf{L})$. $\mathbf{A}=[\bm{a}_{1},\bm{a}_{2},\cdots,\bm{a}_{n}]^{\top}\in\mathbb{R}^{n\times h}$ denotes the atom feature matrix for $n$ atoms within a unit cell, where $\bm{a}_{i}\in\mathbb{R}^{h}$ represents the $h$-dimensional features of atom $i$ in the unit cell. $\textbf{P}=[\bm{p}_{1},\bm{p}_{2},\cdots,\bm{p}_{n}]^{\top}\in\mathbb{R}^{n\times 3}$ is the position matrix, where $\bm{p}_{i}\in\mathbb{R}^{3}$ is the Cartesian coordinate of atom $i$ inside the unit cell. In crystallography, a lattice is a periodic arrangement of points that defines a crystal’s translational symmetry. As shown in Figure˜1, the lattice matrix $\textbf{L}=[\bm{\ell}_{1},\bm{\ell}_{2},\bm{\ell}_{3}]^{\top}\in\mathbb{R}^{3\times 3}$ consists of three translation lattice vectors that determine the shape and periodicity of the unit cell. The periodic repetition of the unit cell is described by integer multiples of the lattice vectors, which establishes the crystal’s inherent translational symmetry. Specifically, the infinite crystal structure can be expressed:
$\hat{\mathbf{P}}=\{\hat{\bm{p}_{i}}|\hat{\bm{p}_{i}}=\bm{p}_{i}+n_{1}\bm{\ell}_{1}+n_{2}\bm{\ell}_{2}+n_{3}\bm{\ell}_{3},n_{1},n_{2},n_{3}\in\mathbb{Z},i\in\mathbb{Z},1\leq i\leq n\}$ (1)
where the integers $n_{1},n_{2},n_{3}$ define a 3D translation with $\bm{\ell}_{1},\bm{\ell}_{2},\bm{\ell}_{3}$. $\hat{\mathbf{A}}=\{\hat{\bm{a}_{i}}|\hat{\bm{a}_{i}}=\bm{a}_{i},i\in\mathbb{Z},1\leq i\leq n\}$ is the atom feature in repeated unit cells, which remains unchanged under periodic translation. In our framework, Cartesian coordinates are used for real space graph construction, where accurate interatomic distances are needed for radius-based cutoff and edge feature initialization.
@@
Fractional coordinates capture the relative positions of atoms within the unit cell and offer significant advantages for periodic materials compared to Cartesian coordinates. Specifically, the fractional coordinate uses the lattice matrix $\textbf{L}\in\mathbb{R}^{3\times 3}$ as the basis, where atomic positions are described by fractional coordinate vectors $\bm{f_{i}}=[f_{1},f_{2},f_{3}]^{\top}\in[0,1)^{3}$. The corresponding Cartesian coordinates are obtained as $\bm{p}_{i}=\bm{L}\bm{f_{i}}$, where $\bm{p}_{i}=f_{1}\bm{\ell}_{1}+f_{2}\bm{\ell}_{2}+f_{3}\bm{\ell}_{3}$. The representation of crystal $\mathbf{M}$ generalizes to $\mathbf{M}=(\mathbf{A},\mathbf{F},\mathbf{L})$, where $\mathbf{F}=[\bm{f}_{1},\cdots,\bm{f}_{n}]^{\top}\in[0,1)^{n\times 3}$ denotes the fractional coordinates of all atoms in the unit cell. Fractional coordinates inherently align with lattice periodicity, ensuring that calculations respect periodic boundary conditions. Moreover, they intrinsically maintain structural relationships under space group operations, such as translations and rotations, consistently preserving the crystal’s underlying symmetry. These properties establish fractional coordinates as a more efficient and symmetry-aware representation for periodic materials compared to Cartesian coordinates, making them essential for the reciprocal space modeling.
@@ 3.2 Reciprocal Space in Crystalline Materials
The periodic arrangement of atoms in crystalline materials fundamentally influences their physical properties by establishing a repeating structure that extends over long distances. Reciprocal space provides a mathematical framework to describe the periodic structure of crystals, enabling the analysis and prediction of material properties influenced by periodicity (Ashcroft & Mermin, 2001; Gross & Marx, 2014; Economou, 2010). Reciprocal space consists of wave vectors $\bm{k}$ spanned by reciprocal lattice vectors $\bm{b_{1}},\bm{b_{2}},\bm{b_{3}}$, which are derived from lattice vectors in real space $\bm{\ell}_{1},\bm{\ell}_{2},\bm{\ell}_{3}$, with relationship satisfying $\bm{b_{i}}\cdot\bm{\ell_{j}}=2\pi\delta_{ij}$, where $\delta_{ij}$ is the Kronecker delta. Each vector in reciprocal space represents a spatial frequency that encodes crystal periodicity along specific directions. More details can be found in Appendix˜A.
To establish the reciprocal space representation, we can use the Fourier transform, which expresses the periodic function $f$ via Fourier series expansion:
@@
where $\hat{f}(\bm{k})$ represents the Fourier coefficients, encoding the contributions of wave vectors $\bm{k}$ at atom $j$. The corresponding inverse transform, which computes $\hat{f}(\bm{k})$ from the real space representation $f(\bm{p}_{j})$, is given by:
@@ 4 Method
Although recent methods have made progress in modeling long-range interactions for crystal property prediction, they often rely on fixed physical formulations (Lin et al., 2023; Taniai et al., 2024) or predefined supercell grids (Kosmala et al., 2023), which limit their adaptability and may break crystal symmetries. In this work, we propose ReciNet, a novel architecture that overcomes these limitations by introducing a learnable reciprocal space block to model long-range periodic interactions. Our proposed ReciprocalBlock, which operates directly on fractional coordinates and reciprocal lattice vectors, learns continuous Fourier-space filters that preserve periodicity and space group symmetries without requiring supercell construction or fixed analytic functions. It operates as a hybrid architecture, synergistically integrating a geometric GNN for short-range effects with the ReciprocalBlock for long-range interactions. For more efficient property prediction, we introduce ReciNet-MT, a multi-task variant that incorporates a task-aware mixture-of-experts (MoE) decoder. The overall architectures of ReciNet and ReciNet-MT are illustrated in Figure˜2.
Figure 2: Overall framework architecture, consisting of three components: (a) Model architecture, (b) ReciNet block, (c) Decoder. The model uses stacked blocks, which integrate local features via message passing and global features via reciprocal space convolution. The decoder outputs property predictions for single-task and multi-task settings.